A force of magnitude 200 N is directed West, and a second force of magnitude 100 N is directed East of
North. Find the magnitude and direction of a third force that, when added to the first two forces, gives a
resultant whose magnitude is zero.
1
Expert's answer
2020-10-30T13:17:40-0400

Answer

"F_1=200N\\F_2=100N"

Assuming that F2 is 30° east of North

resulant force in x -direction

"F_x=200-100sin30°\\=150N"

Force in vertical direction

"F_y=-100cos30\u00b0=-86.6N"

Resulatant force

"F_R=\sqrt{F_x^2+F_y^2}"


"F_R=\sqrt{150^2+86.6^2}=173.2N"

Angle

"tan\theta=\frac{-86.6}{150}\\\theta=-30°"

So third force of 173.2N should be applied -30° below horizontal or 60° east of South.